4x-3y=11
3x+y=18
Solve using multiplication and addition
Help Please!

Answer by Cromlix(4381)

Hi there,
4x-3y=11....(1)
3x+y=18 ....(2)
Multiply (1) by 3
and (2) by 4
12x - 9y = 33....(1)
12x + 4y = 72....(2)
Subtract (2) from (1)
.....-13y = -39
......13y = 39
........y = 3
Substitute y = 3 into
equation (1)
4x - 3y = 11
4x - 3(3) = 11
4x - 9 = 11
4x = 11 + 9
4x = 20
x = 5
Hope this helps:-)